The reaction A→BA→B has been experimentally determined to be second order. The initial rate is 0.0100 M/sM/s at an initial concentration of AA of 0.100 MM.

Part A

What is the initial rate at [A]=0.400M[A]=0.400M?

What is the initial rate at ?

0.00250 M/sM/s
0.160 M/sM/s
0.0100 M/sM/s
0.0400 M/s
Answer #1

Since the reaction is 2nd order in A,

Rate law is:

Rate = k [A]^2

So,

Rate2/rate1 = [A]2^2 / [A]1^2

Rate2/0.0100 = (0.400/0.100)^2

Rate2/0.0100 = (4)^2

Rate2/0.0100 = 16

Rate2= 0.160 M/s

Answer: 0.1600 M/s